40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 22 20:31:58 2010, kiwamu, Update, Green Locking, searching for unknown loss in green PD path RFPDresponse.png
    Reply  Tue Nov 23 01:45:47 2010, kiwamu, Update, Green Locking, noise curve of wideband RFPD noise_RFPD.png
Message ID: 3970     Entry time: Mon Nov 22 20:31:58 2010     Reply to this: 3972
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: searching for unknown loss in green PD path 

 As I said in the past entry (see this entry), there was unknown loss of about 20dB in the beat detection path.

So I started fully characterizing the beat detection path. 

Today I measured the frequency response of the wideband RFPD using the Jenne Laser.

Since all the data were taken by using a 1064nm laser, the absolute magnitudes [V/W] for 532nm are not calibrated yet.  

I will calibrate the absolute values with a green laser which has a known power.


 RFPDresponse.png

The data were taken by changing the bias voltage from -150V to 0V.

The shape of the transfer function looks quite similar to that Hartmut measured before (see the entry).

It has 100MHz bandwidth when the bias voltage is -150V, which is our normal operation point.

 

Theoretically the transfer function must keep flat at lower frequency down to DC.

Therefore for the calibration of this data, we can use the DC signal when a green beam with a known power is illuminating the PD.

 

ELOG V3.1.3-